OpenStudy (anonymous):

4 + 11x = -13 answer as a fraction

OpenStudy (rockstar349002):

1. Subtract four on both sides to get 11x=-17 2. Now divide by 11 on both sides to get x=-17/11 x=-17/11 is your answer Do you understand what I did?

OpenStudy (anonymous):

yes